OpenStudy (anonymous):

what are the x-intercepts of y=-2(x-7)(x+2)?

OpenStudy (nincompoop):

you might want to solve -2(x-7)(x+2) first then set x = 0 to solve for y

OpenStudy (anonymous):

and then?

OpenStudy (nincompoop):

then whatever the answer will be your y-intercept y-intercept means the value(s) that inherently touches the y-axis because x=0

OpenStudy (shamil98):

isn't she asking for the x-intercepts nin

OpenStudy (anonymous):

how do you set it to zero?

OpenStudy (nincompoop):

not y = 0 y-intercept is when x= 0 shamil ohhhhhh nvm

OpenStudy (nincompoop):

I thought I read y-intercept <.<

OpenStudy (shamil98):

-2(x-7)(x+2) = y set y = 0 -2(x-7)(x+2) = 0 and solve for x, to get ur intercepts

OpenStudy (nincompoop):

let me rephrase then

OpenStudy (anonymous):

the answers are -7 and 2 7 and -2 14 and -4 14 and -2

OpenStudy (anonymous):

so i dont think you set it to zero

OpenStudy (nincompoop):

so if you're looking for the x-intercept then set your y to zero then solve since your equation is already in factored form, solve it like this x-7 = 0 and x+2 = 0, the -2 outside is automatically zero

OpenStudy (anonymous):

so then its 7, -2 you do opposite the signs?

OpenStudy (nincompoop):

x-7 = 0 x - 7 + 7 = 0 + 7 x= ? x+2 = 0 x+ 2 + (-2) = 0 + (-2) x = ?

OpenStudy (shamil98):

-2(x-7)(x+2) = 0 divide by -2 to both sides (x-7)(x+2) = 0 x = 7 , x= -2

OpenStudy (anonymous):

ooo I get it thank you!